OpenStudy (anonymous):

Find The Equation of the tangent to the curve : x^2+y^2=25 at x=3

OpenStudy (anonymous):

Come on, what's the problem with this one?

OpenStudy (anonymous):

lol,dude ..i want only to check the answers of these questions ,i know their answers all :)

OpenStudy (experimentx):

m = -x/y x=3, y=+-4 use slope point formula you will have two tangents

OpenStudy (anonymous):

he want the equation ,and i think it has only one equation for one tangent !!!

OpenStudy (experimentx):



OpenStudy (anonymous):

that at x =3 only ,right ?

OpenStudy (anonymous):

agreed but usually we're only concerned with the positive half of the circle so we can still remain under the pretense that this is operating as a true function (i.e. passes the verticle line test)

OpenStudy (anonymous):

ha i just realized that I posted the integral instead of the derivative, my bad guys. take the derivative in terms of x and then plug 3 in for x and you can solve

OpenStudy (experimentx):

use implicit differentiation, \[ 2x + 2y\frac{dy}{dx} = 0\] \[ \frac{dy}{dx} = -x/y\] find the value of y from plugging the value of x, use this to find slope ... and rest is line stuff.

OpenStudy (anonymous):

aha ,ty :)

OpenStudy (experimentx):

wc